OpenStudy (anonymous):

PLEASE HELP. Find the standard form of the equation of the parabola with a focus at (7, 0) and a directrix at x = -7 a. y 1/28 = x^2 b. x 1/28 = y^2 c. -28y = x^2

OpenStudy (anonymous):

conic sections

OpenStudy (anonymous):

find the vertex

OpenStudy (anonymous):

is there a picture

OpenStudy (anonymous):

There is no picture given.

OpenStudy (tkhunny):

Focus-directrix definition is a wonderful thing. Locus of points Equidistant from a given focus and a given directrix. How far are points from (7,0)? \(\sqrt{(x-7)^{2}+y^{2}}\) How far are points from x = -7? \(|x-(-7)| = |x+7|\) That's all you need. Just a little algebra to go!

OpenStudy (tkhunny):

You can also just jump right to the vertex. (7,0) and x = -7 says clearly that the vertex is (0,0). Just draw in the axis of symmetry and you will see it.

OpenStudy (anonymous):

So it'd be x 1/28 = y^2 because it's on the x-axis?

OpenStudy (anonymous):

Wait I mean y 1/28 = x^2

OpenStudy (anonymous):

right choice is not on the list

OpenStudy (anonymous):

i think it's\[\sqrt{(x-7)^{2}+(y-0)^{2}} = \sqrt{(y+7)^{2}}\]

OpenStudy (anonymous):

y2 = 14x?

OpenStudy (anonymous):

Is that right? @sourwing

OpenStudy (anonymous):

now a little algebra is used

OpenStudy (anonymous):

How so? @faisalalif1999

OpenStudy (anonymous):

one xec

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

y^2 = 28x

OpenStudy (cwrw238):

I showed you how to do this already

OpenStudy (anonymous):

Haha sorry, I wasn't convinced that you were right. Thanks.

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

cwrw238 is right

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

What is a directrix?

OpenStudy (cwrw238):

http://www.mathwords.com/d/directrix_parabola.htm - will tell you about the directrix

OpenStudy (anonymous):

Thanks a ton! I appreciate this!